Direct Materials and Direct Labor Variance Analysis

Abbeville Fixture Company manufactures units in a small manufacturing facility. The units are made from brass. Manufacturing has 30 employees. Each employee presently provides 35 hours of labor per week. Information about a production week is as follows:

Standard wage per hour

$16.2

Standard labor time per unit

10 min.

Standard number of lbs. of brass

2.1 lbs.

Standard price per lb. of brass

$12

Actual price per lb. of brass

$12.25

Actual lbs. of brass used during the week

18,169 lbs.

Number of units produced during the week

8,400

Actual wage per hour

$16.69

Actual hours for the week (30 employees × 35 hours)

1,050 hrs.

Required:

a.  Determine the standard cost per unit for direct materials and direct labor. Round the cost per unit to two decimal places.

Direct materials standard cost per unit

$

Direct labor standard cost per unit

$

Total standard cost per unit

$

b.  Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Round your answers to the nearest whole dollar. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Materials Price Variance

$

Direct Materials Quantity Variance

$

Total Direct Materials Cost Variance

$

c.  Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Round your answers to the nearest whole dollar. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Labor Rate Variance

$

Direct Labor Time Variance

$

Total Direct Labor Cost Variance

$
Answer #1
a
Direct materials standard cost per unit 25.20 =2.1*12
Direct labor standard cost per unit 2.70 =16.2*(10/60)
Total standard cost per unit 27.90
b
Direct Materials Price Variance 4542 Unfavorable =18169*(12.25-12)
Direct Materials Quantity Variance 6348 Unfavorable =12*(18169-8400*2.1)
Total Direct Materials Cost Variance 10890 Unfavorable =(18169*12.25)-(8400*2.1*12)
c
Direct Labor Rate Variance 515 Unfavorable =1050*(16.69-16.2)
Direct Labor Time Variance -5670 Favorable =16.2*(1050-8400*10/60)
Total Direct Labor Cost Variance -5156 or -5155 Favorable =(1050*16.69)-(8400*16.2*10/60)